# STG-MTL: Scalable Task Grouping for Multi-Task Learning Using Data Maps

### v1

@@ 4.1 Datasets and Tasks
Our task generation is based on the CIFAR10 and CIFAR100 datasets (Krizhevsky et al., 2009). We define three groups of tasks for our experiments. In Group 1 (G1), we include binary classification tasks that determine whether an image belongs to a specific label in CIFAR10 or not. G1 consists of 10 tasks: airplane, automobile, bird, cat, deer, dog, frog, horse, ship, truck.
Group 2 (G2) expands on G1 by introducing additional tasks on CIFAR10. These tasks include Living being, Odd-numbered, Downside, Not living being, random. The “Living being” task aims to detect whether an image contains a living being, which includes images with the labels bird, cat, deer, dog, frog, horse. Similarly, the “Not living being” task focuses on identifying non-living beings; these are airplane, automobile, ship, truck classes in CIFAR10. Notably, “Living being” and “Not living being” are intentionally designed to be similar tasks for testing purposes. The “Odd-numbered” task identifies whether the label of a CIFAR10 image is odd or not, encompassing automobile, cat, dog, horse, truck classes. Additionally, we flip half of CIFAR10 images and create a task to train the model to recognize vertically flipped images, the “Downside” task. Lastly, the “random” task assigns random binary labels to the entire dataset with a predefined seed for consistency and reproducibility. It is worth mentioning that while the original tasks in G1 are imbalanced, the extra tasks in G2 are all balanced ones.
Group 3 (G3), similar to G1, consists of 100 binary classification tasks using the CIFAR100 labels. We also utilize the super labels of CIFAR100 as our ground truth for task clustering evaluation. It is worth mentioning that CIFAR100 super labels are not intended for task grouping, so they are not grouped based on visual similarities like our method’s objective. Instead, they are mostly clustered semantically, even though there are some exceptions like mushrooms and the classes of vehicles 1 and 2. Still, we think they serve as an informative indicator of the effectiveness of our approach, especially in the visually coherent superclasses.
@@ 4.2 Model Architectures and Hyper-Parameters
For all our experiments, we adopt the RESNET18 architecture (He et al., 2016) as our base model. Our method is model-agnostic, so we have also experimented with models with much less complexity, yet we use RESNET18 considering it has moderate model capacity. Furthermore, we utilize it without any pre-training, ensuring that the model starts from scratch for each task grouping scenario. The last fully connected layer of RESNET18 serves as the task heads, with the number of output neurons corresponding to the number of tasks. Each neuron in the task heads represents a specific classification task. Throughout our experiments, the rest of the network, excluding the task heads, is shared among all tasks. Also, we train the model for epochs in all our experiments: to extract data maps and to evaluate the models. Additionally, in our clustering process, we primarily set the fuzzification index () to , unless explicitly mentioned otherwise. The fuzzification index controls the level of fuzziness in the soft clustering algorithm, so increasing it produce softer decisions.
@@ 4.3 Task Clustering Results
As illustrated in Figure 5, even without additional tasks, our method performed the same reasonable clustering for G1, grouping living beings together and non-living things together 5. Additionally, Figure 5 demonstrates the clustering using three clusters, revealing that the living being cluster was divided into two groups: cluster and cluster . Cluster predominantly contained quadruped animals cat, dog, horse, while cluster included bird, frog, deer that represented the other living creatures except for the deer. These results showcase the effectiveness of our clustering algorithm in capturing semantic similarities among tasks based on the visual data leading to meaningful task groupings.
In addition to our experiments on G1 and G2, we conducted a comprehensive evaluation of an unprecedented number of tasks, specifically tasks from CIFAR100, in G3. As part of this evaluation, we compared our task clustering results against the predefined superclasses provided by CIFAR100. It is important to note that the superclasses in CIFAR100 primarily rely on semantic relationships as illustrated in Section 4.1. That is why we focus on coherent superclasses like people and flowers, as examples.
In Figure 6, we showcase an example of the clustering results for a group of super tasks. It is noteworthy that our method successfully clusters certain groups of tasks in alignment with the predefined CIFAR100 superclasses, as illustrated in Figure 6. However, it is important to acknowledge that there are cases where the clustering may not be perfect, as depicted in Figure 6; we think this is primarily because our method focus one visual similarities, which is exploited during training rather than semantics. Nevertheless, even in such instances, our clustering algorithm manages to allocate significant weights of all tasks into distinctive clusters, such as clusters and in Figure 6. Notably, in cluster , the participation percentages of the tasks orchid, poppy, rose, tulip are the highest across all clusters, indicating a close relationship with the missclassified task sunflower, yet our method suggests that the other four tasks are more visually related. We further discuss all the clustering result details of the 100 Tasks in Appendix A.
@@ 4.4 Evaluation Analysis
To further validate the effectiveness of our method, we conducted a comprehensive evaluation as described in Section 3.4 on all task groups. Figure 7 presents the average F1 score for both the training and test sets of all the three sets of tasks. Our method is denoted by STG-MTL xxC (F=2) where xx represents the number of clusters. The MTL curve represents the results obtained from training an MTL model on all tasks without any grouping, while the STL curve represents the results obtained by training separate STL models for each task and merging their outputs. We compare the performance of our method against the MTL and STL approaches in both G1 & G2 and against the MTL approach only in G3 because the STL performance is poorer than the MTL, as it overfits.
Overall, our method consistently outperforms both the MTL and STL approaches, indicating that the task grouping provides valuable information for improving task performance. Notably, although our method tends to overfit and achieves excellent training performance, it also achieves the best performance on the test set. This suggests that if the models were further fine-tuned, even greater gains could be achieved, yet we refrain from tuning any of the models in this study to guarantee fairness in comparison.
@@ 5 Conclusion and Future Work
In conclusion, we have presented STG-MTL, which is a novel scalable approach for task grouping in multi-task learning (MTL) settings. Our method utilizes data maps (Swayamdipta et al., 2020) to identify task similarities and group them accordingly. We showed is superior scalability theoretically in comparison to TAG (Fifty et al., 2021), HOA (Standley et al., 2020), and MTG-Net (Song et al., 2022). We have also demonstrated the effectiveness of our method through our experiments on CIFAR10 and CIFAR100 datasets (Krizhevsky et al., 2009), where we pushed the boundaries by experimenting with 100 tasks, which has never been done before in the literature proving its scalability. We have also compared our clustering results against the predefined superclasses in CIFAR100, further validating the effectiveness of our approach. Furthermore, our method outperformed traditional MTL and single-task learning (STL) approaches, showcasing the quality of task grouping and its ability to improve multi-task learning performance.
For future work, we plan to expand the scope of our experiments by including a wider range of datasets and task types, enabling a more comprehensive evaluation of our approach’s effectiveness and applicability. Furthermore, as our Data Maps are currently limited to classification tasks, we aim to explore their generalization to other task types, such as regression. Additionally, we hope our research could open a new research direction in the MTL community to explore the development of new features that can capture the training dynamics efficiently, other than data maps. By advancing this research direction, we can unlock new possibilities for enhancing performance and driving further advancements in the field of MTL.
Acknowledgments We gratefully thank the Fatima Fellowship111https://www.fatimafellowship.com/ for supporting this research especially during its early stages.
@@ Appendix A Task Clustering Results of 100 Tasks
Notably, the clustering with clusters demonstrates successful grouping in many categories, such as People, Trees, Food Containers, Flowers, Household Electrical Devices. Figures 8 and 8 highlight the close association between Vehicles 1 and 2, as they are almost merged into the same cluster (Cluster 15). When reducing the number of clusters to , we observe enhanced coherence in the assigned tasks. For instance, the tasks related to ”Fruit and Vegetables” are nearly clustered together after reducing the number of clusters.
Moreover, our method effectively captures the logical associations between categories. Figures 9 and 9 showcase the plausible merge between ”Household Electrical Devices” and ”Household Furniture.” Additionally, we observe similar distribution patterns among related categories, such as Large Carnivores, Large Omnivores and Herbivores, Medium-Sized Mammals in Figures 9, 9, and 9 respectively.

### v2

@@ 4.1 Datasets and Tasks
Our task generation is based on the CIFAR10 and CIFAR100 datasets [17] along with CelebA dataset [21]. We mainly use them due to their public availability and the available high number of tasks we could generate from them. We show the results of four groups of tasks in our experiments. For all of these datasets, we define binary tasks to detect whether the input image belongs to a particular label or not. In Group 1 (G1), for example, we include binary classification tasks that determine whether an image belongs to CIFAR10 labels. G1 consists of 10 tasks: airplane, automobile, bird, cat, deer, dog, frog, horse, ship, truck. That is the “airplane” task is to detect whether the image is an airplane or not.
Group 2 (G2) expands on G1 by introducing additional tasks on CIFAR10, designed for testing qualitatively the method behavior. These tasks include Living being, Odd-numbered, Downside, Not living being, random. The “Living being” task aims to detect whether an image contains a living being, which includes images with the labels bird, cat, deer, dog, frog, horse. Similarly, the “Not living being” task focuses on identifying non-living beings; these are airplane, automobile, ship, truck classes in CIFAR10. Notably, “Living being” and “Not living being” are intentionally designed to be similar tasks to check if the method can detect such similarity. The “Odd-numbered” task identifies whether the label of a CIFAR10 image is odd or not, encompassing automobile, cat, dog, horse, truck classes. Additionally, we flip half of CIFAR10 images and create a task to train the model to recognize vertically flipped images, the “Downside” task. Lastly, the “random” task assigns random binary labels to the entire dataset with a predefined seed for consistency and reproducibility; the “random” task is added to represent extremely difficult task, for example. It is worth mentioning that while the original tasks in G1 are imbalanced, the extra tasks in G2 are balanced.
Group 3 (G3), similar to G1, consists of binary classification tasks using the CIFAR100 labels. We also utilize the super labels of CIFAR100 as our ground truth for task clustering evaluation. It is worth mentioning that CIFAR100 super labels are not intended for task grouping, so they are not grouped based on visual similarities like our method’s objective. Instead, they are mostly clustered semantically, even though there are some exceptions like mushrooms and the classes of vehicles 1 and 2. Still, we think they serve as an informative indicator of the effectiveness of our approach, especially in the visually coherent superclasses. Similarly, we define group 4 (G4) to include the binary classification tasks included in CelebA.
@@ 4.2 Model Architectures and Hyper-Parameters
For all our experiments, we adopt the RESNET18 architecture [10] as our base model. Our method is model-agnostic, so we have also experimented with simpler models, results & archtecture discussed in Appendix A.1, yet we use RESNET18 considering its moderate model capacity. Furthermore, we utilize it without any pre-training, ensuring that the model starts from scratch for each task grouping scenario. The last fully connected layer of RESNET18 serves as the task heads, with the number of output neurons corresponding to the number of tasks. Each neuron in the task heads represents a specific classification task. Throughout our experiments, the rest of the network, excluding the task heads, is shared among all tasks. Also, we train the model for epochs in all our experiments: to extract data maps and to evaluate the models. Additionally, in our clustering process, we primarily set the fuzzification index () to . The fuzzification index controls the level of fuzziness in the soft clustering algorithm, so increasing it produce softer decisions.
@@ 4.3 Task Clustering Results
As illustrated in Figure 4, even without additional tasks, our method performed the same reasonable clustering for G1, grouping living beings together and non-living things together. Additionally, Figure 4 demonstrates the clustering using three clusters, revealing that the living being cluster was divided into two groups: cluster and cluster . Cluster predominantly contained quadruped animals cat, dog, horse, while cluster included bird, frog, deer that represented the other living creatures except for the deer. These results showcase the effectiveness of our clustering algorithm in capturing semantic, visual similarities among tasks based on the visual data leading to meaningful task groupings.
In addition to our experiments on G1 and G2, we conducted a comprehensive evaluation of an unprecedented number of tasks, specifically tasks from CIFAR100, in G3. As part of this evaluation, we compared our task clustering results against the predefined superclasses provided by CIFAR100. It is important to note that the superclasses in CIFAR100 primarily rely on semantic relationships as illustrated in Section 4.1, so we show the results of people and flowers, as examples.
In Figure 5, we showcase an example of the clustering results for a group of super tasks. It is noteworthy that our method successfully clusters certain groups of tasks in alignment with the predefined CIFAR100 superclasses, as illustrated in Figure 5. However, it is important to acknowledge that there are cases where the clustering may not be perfect, as depicted in Figure 5; we think this is primarily because our method focus one visual similarities, which is exploited during training rather than semantics. Nevertheless, even in such instances, our clustering algorithm manages to allocate significant weights of all tasks into distinctive clusters, such as clusters and in Figure 5. Notably, in cluster , the participation percentages of the tasks orchid, poppy, rose, tulip are the highest across all clusters, indicating a close relationship with the missclassified task sunflower, yet our method suggests that the other four tasks are more visually related. We further discuss all the clustering results of the 100 Tasks in Appendix A for RESNET18; we also show the clustering results of simpler CNN model in Appendix A.1, which shows the possible potential of transferring the results generated from simpler models to other models.
@@ 4.4 Evaluation Analysis
To further validate the effectiveness of our method, we conducted a comprehensive evaluation as described in Section 3.4 on all task groups. Figure 6 presents the average F1 score for both the training and test sets of all the three sets of tasks; we use F1 as our evaluation metric becuase many tasks are imbalanced, so we use F1 score to adhere to such issue. Our method is denoted by STG-MTL xxC (F=2) where xx represents the number of clusters. The MTL curve represents the results obtained from training an MTL model on all tasks without any grouping, while the STL curve represents the results obtained by training separate STL models for each task and merging their outputs. We compare the performance of our method against the MTL and STL approaches in both G1 & G2 and against the MTL approach only in G3 & G4 because the STL performance is poorer than the MTL, as it overfits.
We also present additional experiments, figure 7 using a simpler model with much less number of parameters, of a custom CNN architecture depicted in Table 2, in Appendix A.1. These experiments demonstrate the robustness and adaptability of our method across different model architectures. It is worth mentioning that even with hard clustering, our method outperforms standard MTL, showcasing its effectiveness in leveraging task grouping for enhanced performance as shown in Figures 7 and 7. Moreover, the results obtained using our approach with data maps extracted from the MTL model are comparable to those from the STL models as in Figures 7 and 7.
Overall, our method consistently outperforms both the MTL and STL approaches, indicating that the task grouping provides valuable information for improving task performance. Notably, although our method tends to overfit, in RESNET18 experiments, and achieves excellent training performance, it also achieves the best performance on the test set. This suggests that if the models were further fine-tuned, even greater gains could be achieved, yet we refrain from tuning any of the models in this study to guarantee fairness in comparison.
@@ 5 Conclusion and Future Work
In conclusion, we have presented STG-MTL, which is a novel scalable approach for task grouping in multi-task learning (MTL) settings. Our method utilizes data maps [33] to identify task similarities and group them accordingly. We showed its superior scalability theoretically in comparison to TAG [9], HOA [30], and MTG-Net [29]. We have also demonstrated the effectiveness of our method through our experiments on CIFAR10 & CIFAR100 [17] and CelebA [21] datasets , where we pushed the boundaries by experimenting with 100 tasks, which has never been done before in the literature proving its scalability. We have also compared our clustering results against the predefined superclasses in CIFAR100, further validating the effectiveness of our approach. Nevertheless due to our limited computational power and the poor scalability of the current methods, we were not able to evaluate the performance of our results against the other methods. Instead, we showed that our method outperformed both traditional MTL and single-task learning (STL) approaches, showcasing the quality of task grouping and its ability to improve multi-task learning performance.
For future work, we aim to explore Data Maps’ generalization to other task types, such as regression, because they are currently limited to classification tasks only. Additionally, we hope our research could open a new research direction in the MTL community to explore the development of new features that can capture the training dynamics efficiently, other than data maps. By advancing this research direction, we can unlock new possibilities for enhancing performance and driving further advancements in the field of MTL.
Acknowledgments We gratefully thank the Fatima Fellowship222https://www.fatimafellowship.com/ for supporting this research especially during its early stages.
@@ Appendix A Task Clustering Results of 100 Tasks
Notably, the clustering with clusters demonstrates successful grouping in many categories, such as People, Trees, Food Containers, Flowers, Household Electrical Devices. Figures 8 and 8 highlight the close association between Vehicles 1 and 2, as they are almost merged into the same cluster (Cluster 15). When reducing the number of clusters to , we observe enhanced coherence in the assigned tasks. For instance, the tasks related to “Fruit and Vegetables” are nearly clustered together after reducing the number of clusters.
Moreover, our method effectively captures the logical associations between categories. Figures 9 and 9 showcase the plausible merge between “Household Electrical Devices” and “Household Furniture.” Additionally, we observe similar distribution patterns among related categories, such as Large Carnivores, Large Omnivores and Herbivores, Medium-Sized Mammals in Figures 9, 9, and 9 respectively.
@@
A.1 Robust Clustering using Less Complex Architectures
Layer | Kernel Size | Output |
Convolution | channels | |
2D BN + ReLU | ||
2D MaxPool | ||
Convolution | channels | |
2D BN + ReLU | ||
Convolution | channels | |
2D BN + ReLU | ||
2D MaxPool | ||
Flatten | ||
FC | ||
FC |
To assess the robustness and generalization potential of our approach, we conducted a similar experiment using a simpler CNN architecture. The architecture’s details are provided in Table 2, where . Notably, the CNN architecture used for extracting the data maps in our clustering process comprises only learnable parameters. In contrast, the ResNet18 architecture, from which we initially derived our clustering results, consists of learnable parameters.
Figure 10 and 11 showcase the task clustering results obtained using the simplified CNN architecture. Surprisingly, the resulting clustering patterns from our custom CNN are remarkably comparable to those achieved with ResNet18. This observation underscores the method’s expressiveness, as it effectively captures task relationships without necessitating a high number of learnable parameters.
This intriguing finding opens the possibility of extrapolating clustering outcomes from less expressive models to more complex ones. However, further investigations are required to ascertain the extent and implications of this approach.
Appendix B Modularity in the Code Structure
In designing our method, we placed significant emphasis on scalability and modularity. Ensuring scalability without requiring substantial modifications whenever model architecture or tasks change, code modularity, is crucial. To achieve this, we designed a modular pipeline that facilitates easy adaptation to various scenarios with minimal adjustments. The class diagram for our pipeline is depicted in Figure 12.
Our code architecture is composed of several classes, each with distinct roles. The main class responsible for training models on datasets is the Trainer. Notably, the Trainer employs an object from the DataMapRecorder class, which dynamically updates and records data maps after each iteration. To enhance reusability, we introduce an abstract class called BaseModel, which serves as the blueprint for all models. Similarly, the Trainer expects a BaseDataModule as input, an abstract class encompassing dataset information and instructions for building data loaders.
When dealing with new sets of tasks, two steps are necessary. First, we inherit from BaseModel, implementing its abstract methods, italicized in the figure. Second, we provide concrete implementations for the BaseDataModule’s abstract methods. Our code contains examples of these two steps for different datasets such as CIFAR10, CIFAR100, and CelebA.
To train models and extract data maps, only a few lines of code are required, exemplified in Code 1. The resulting data map from training an MTL model is then ready for clustering using our DatamapTaskClustering class, depicted in Figure 13, as shown in Code 2. This design simplifies the integration process for new tasks, models, or datasets, effectively enhancing the overall modularity of our pipeline and method.